Is 3/x sometimes, always, or never an integer
1
Expert's answer
2012-09-18T10:39:26-0400
The function f(x) = 3/x "sometimes" takes integer values, that is it also takes
non-integer values.
Actually f(x) is integer if and only if x = 3/n, where n
is a non-zero integer.
In this case f(3/n) = 3/(3/n) = n.